# Hyperparameter Optimization in Machine Learning

### v2

@@ 4.1 Model-based HPO: a general schema
2:Warm-up dataset $\mathcal{C}_{0}=\{(\lambda_{j},f(\lambda_{j}))\}$
@@
5:$\beta\leftarrow\min_{j}f(\lambda_{j})$triangleright Initialize incumbent
6:$\mathcal{C}_{1}\leftarrow\mathcal{C}_{0}$
@@
8:     Fit $\hat{f}_{k}$ to $\mathcal{C}_{k}$triangleright Update step ($\pi_{2}$)
9:     $\lambda_{k}\leftarrow\operatorname*{arg\,max}_{\lambda^{\prime}}\alpha(\lambda^{\prime}|\hat{f}_{k-1},\beta)$triangleright Decision step ($\pi_{1}$)
10:     Evaluate $f(\lambda_{k})$triangleright Expensive step
11:     if $f(\lambda_{k})\leq\beta)$ then
12:         $\beta\leftarrow f(\lambda_{k})$
13:     end if
14:     $\mathcal{C}_{k+1}\leftarrow\mathcal{C}_{k}\cup(\lambda,f(\lambda))$
@@
The vast majority of model-based methods is, in principle, sequential: the surrogate model suggests the next point to evaluate while being simultaneously adapted online as new data from the execution of the learning algorithm $A$ is collected and processed.
It can be useful to think about a MB-HPO method as an iterative process over both the hyperparameter space $\Lambda$ and the surrogate function space $\hat{\mathcal{F}}$. To represent this process, we introduce two mappings $\pi_{1}:\mathcal{F}\times\mathcal{X}\to\Lambda$ and $\pi_{2}:(\Lambda\times\mathbb{R})^{*}\times\hat{\mathcal{F}}\times\mathcal{X}\to\hat{\mathcal{F}}$, one corresponding to the iteration in the hyperparameter space, the other to the update of the surrogate model. Then, we can describe the iterates of a MB-HPO method as follows:
$\left(\begin{array}[]{c}\lambda_{k+1}\\ \hat{f}_{k+1}\end{array}\right)=\left(\begin{array}[]{c}\pi_{1}(\hat{f}_{k},\upsilon_{k})\\ \pi_{2}\left(\mathcal{C}_{k},\hat{f}_{k},\upsilon_{k}\right)\end{array}\right),$ (4.3)
where the variables $\nu_{k}$ represent auxiliary inputs and $\mathcal{C}_{k}$ a set of paris of configurations and response function evaluations. The surrogate update mapping $\pi_{2}$ may often entail an online update of $\hat{f}$, therefore depending only on the last collected point $(\lambda_{k},f(\lambda_{k}))$ and the current surrogate model $\hat{f}_{k}$.
Typically $\hat{f}_{0}$ is initialized by fitting it to several evaluations obtained in a “warm-up stage”, e.g., where configuration are generated by random search or Latin hypercube sampling. Then, the points $\mathcal{C}_{k}=\mathcal{C}_{0}\cup\{(\lambda_{j},f(\lambda_{j}))\}_{j=1}^{k}$ that comprise past configurations and evaluations of the response function (including warm-up evaluations) can be though of as a dataset to fit the surrogate model. We summarize the typical steps of a model-based method in Algorithm 1, where the decision map $\pi_{i}$ is implemented by the optimization of an AF $\alpha$. The algorithm generates a sequence of $T$ configurations. As in the previous chapter, we denote by $\hat{\lambda}_{k}$ the best configuration found after $k$ steps.
@@ 4.2 Bayesian optimization
BO has long been used as a method to globally optimize black-box functions, that is, functions for which we may only observe the value $f(\lambda)$ at any point $\lambda\in\Lambda)$. It dates back at least to the seventies with a series of works by J. Močkus [mockus:1975:bayesian_methods_seeking, mockus1978application]. While limited to small-to-medium dimensional settings, BO is an effective approach to optimize (highly) multi-modal, non-convex and possibly expensive-to-compute functions. As a result, it has attracted considerable attention for its potential to tackle HPO problems, especially following the publication of three influential papers in the early 2010s [bergstra2011algorithms, hutter_sequential_2011, snoek_practical_2012]. In this section, we give a brief overview of the general methodology in the HPO context. We refer the reader to the work by jones1998efficient for an intuitive introduction of the mathematical aspects of the methodology and to the review by shahriari:2016:taking_human_out for further insights into practical aspects and applications to machine learning.
In Bayesian optimization, $\hat{y}$ is a stochastic process rather than a function and $\hat{y}(\lambda)$ is a random variable. Here randomness arises because of the epistemic uncertainty of the surrogate model (due to the small size of $\mathcal{D}$) about the response at $\lambda$. This surrogate model is fitted to available observations by Bayesian conditioning, so that afterwards it represents the Bayesian posterior distribution, which is used to decide the next value to probe. Such decision is most often expressed as the maximization of an acquisition function, as we introduced before. The AF balances exploration — collecting information to improve the model itself, by proposing points away from previous observations — and exploitation — picking points close to the current minimizer in order to refine the best solution. Thus, referring to Eq. (4.3), a typical BO method implements $\pi_{1}$ with the optimization of an AF and $\pi_{2}$ with a Bayesian update of the stochastic surrogate model.
First, we discuss standard choices to design and fit surrogate models in BO, including Gaussian processes and tree Parzen estimators. Then, we focus on AFs, especially on the expected improvement [mockus1978application]. One may say that a most classic instantiation of BO consists in modelling the surrogate with a Gaussian process and using as AF the expected improvement. This approach is well-suited for the global optimization of smooth multivariate real functions, and it provides an accessible setting to approach BO for hyperparameter optimization. Nevertheless, it may not be the most suitable for some HPO problems; thus we discuss limitations along the way.
@@ 4.2.1 Bayesian linear model
We start by introducing a simple linear model $h_{w}$ in a given feature space $\mathcal{V}$, to which we add independent Gaussian noise with variance $\sigma^{2}$:
@@
where $\varphi:\Lambda\to\mathcal{V}$ is a feature map. While using linear maps directly in the input space would result in an overly simplistic model – we are trying to model an objective, the validation error, which can be (highly) non-convex and multimodal – working in a different feature space allows us to capture much richer behaviors but still retain most of the benefits of the linear formulation [murphy2012machine, Ch. 14]. As evaluating $f$ is expensive, we expect to work with only a few observations. The noise in Eq. (4.4) models potential errors in the observations (see also 4.2.4).
@@
Suppose now we are at the $k$-th iteration of the algorithm, after having already collected $k$ observations $\mathcal{C}=\{(\lambda_{i},f(\lambda_{i}))\}_{i\in[k]}$. Let $Y_{k}=\left(f(\lambda_{i})\right)_{i=1}^{k}\in\mathbb{R}^{k}$ be the vector of observed values, and denote by $X_{k}=\left(\varphi(\lambda_{1}),\dots,\varphi(\lambda_{k})\right)$ the corresponding design matrix in the feature space (organized by columns). We are interested in obtaining a probability distribution over the output of $f$ at a query point $\lambda_{*}\in\Lambda$, that is, in computing
$y_{*}\sim\hat{f}_{k+1}(\lambda_{*})=p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k}),$
where $x_{*}=\varphi(\lambda_{*})$ is a query point in the feature space. This is going to be useful later in Section 4.2.3, where an AF employs $\hat{f}_{k}$ to select the next configuration $\lambda_{k+1}$.
Modelling uncertainty away from observations.
@@
We could follow an empirical risk minimization approach, obtaining the weights $w_{k}\in\mathbb{R}^{\mathrm{dim}(\mathcal{V})}$ of the model Eq. (4.4) by minimizing, for example, a (regularized) mean squared error over the dataset $(X_{k},Y_{k})$. We could then directly set $\hat{f}_{k+1}=h_{w_{k}}$. Such an estimate would, however, leave us with a rather meager predictive distribution $p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w_{k})$, capable of capturing only our belief of the amount of noise present in the observations. Rather, we would like the stochasticity of $\hat{f}$ to (primarily) represent our uncertainty about the behavior of $f$ in unexplored regions. To better grasp the difference between these two aspect, consider the case where $f$ is a (non-linear) smooth deterministic function.  [^1] The following argument is mutated from jones1998efficient. If we were to follow an ERM approach, we could either omit the noise term $\varepsilon$ in Eq. (4.4), or leave it for capturing modelling errors. However, the first scenario, would yield a single point estimate for any $\lambda\in\Lambda$, which is very unlikely to be accurate, especially for points far from those previously observed. Conversely, if we still wish to maintain a noise component, $\hat{f}_{k+1}$ would trivially be inaccurate at $\lambda_{i}$ for $i\leq k$ (as $f$ is deterministic). Thus, both ERM modelling options would leave us with a surrogate model that is almost surely inaccurate.
@@
A way to escape this apparent stalemate is to replace the independent Gaussian noise with one that depends on the evaluation point, i.e. $\varepsilon=\varepsilon(\lambda)$. Then, when fitting the model to the $k$ observations, one may have $\varepsilon_{k}(\lambda_{i})=0$ for $i\leq k$ while letting $\varepsilon_{k}(\lambda)\neq 0$ for $\lambda\neq\lambda_{i}$. Furthermore, by continuity ($\varepsilon_{k}$ is smooth and continuous as difference of $C^{1}$ functions), $\varepsilon_{k}(\lambda)\to\varepsilon_{k}(\lambda_{i})=0$ as $\lambda\to\lambda_{i}$. The Bayesian perspective offers a principled way to achieve this behaviour.
Thus, we can reason about all the linear models (in the feature space) that agree with the observations, weighted by their “realization probability”.
More formally, we marginalize over the weights $w$ and set
$\hat{f}_{k+1}(\lambda_{*})=\int p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w)p_{w}(w|X_{k},Y_{k})\,\mathrm{d}w.$ (4.5)
@@
Given our modeling assumption (4.4) and the hypothesis that the noise is independent, the first distribution simply reads as
$p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k},w)=p_{y_{*}}(\cdot|x_{*},w)=\mathcal{N}(h_{w}(\lambda_{*}),\sigma_{\varepsilon}^{2})=\mathcal{N}(x_{*}^{\intercal}w,\sigma_{\varepsilon}^{2}).$ (4.6)
By the Bayes’ rule, $p_{w}(\cdot|X_{k},Y_{k})$ is instead given by
$p_{w}(\cdot|X_{k},Y_{k})=\frac{p_{y}(\cdot|X_{k},w)p_{w}(\cdot|X_{k})}{p_{y}(\cdot|X_{k})}.$ (4.7)
@@
$p_{y}(\cdot|X_{k},w)=\mathcal{N}(X_{k}^{\intercal}w,\sigma_{\varepsilon}^{2}I).$
@@
$p_{w}(\cdot|X_{k})=p_{w}=\mathcal{N}(0,\Sigma),$ (4.8)
where $\Sigma$ is a positive definite covariance matrix. This is a so-called prior. Equipped with Eq. (4.8), we can now also compute the denominator of (4.7) by marginalizing $p_{y}(\cdot|X_{k})$ over $w$. We obtaining
$p_{y}(\cdot|X_{k})=\int p_{y}(\cdot|X_{k},w)p_{w}(w)\,\mathrm{d}w=\mathcal{N}(0,X_{k}^{\intercal}\Sigma X_{k}+\sigma_{\varepsilon}^{2}I)$ (4.9)
We recognize, now, that the feature map only appears in a quadratic form where the $(i,j)$-th entry is given by $\varphi(\lambda_{i})^{\intercal}\Sigma\,\varphi(\lambda_{j})$: a dot product in the feature space. We then introduce the kernel corresponding to the feature map $\phi$ as follows:
@@
and rewrite Eq. (4.9) in terms of the Gram matrix $K=\{\kappa_{\Sigma}(\lambda_{i},\lambda_{j})\}_{i,j=1}^{k}\in\mathbb{R}^{k\times k}$ as
$p_{y}(\cdot|X_{k})=\mathcal{N}(0,K+\sigma_{\varepsilon}^{2}I).$
We finally have all the terms to compute (4.5); which is, again, Gaussian:
$\hat{f}_{k+1}(\lambda_{*})=p_{y_{*}}(\cdot|x_{*},X_{k},Y_{k})=\\ \mathcal{N}(k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)^{-1}Y_{k},\kappa(\lambda_{*},\lambda_{*})-k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)k_{*})$ (4.11)
where $k_{*}=\{\kappa_{\Sigma}(\lambda_{*},\lambda_{i})\}_{i=1}^{k}\in\mathbb{R}^{k}$ is the vector of kernel evaluations between the query and the observations points. Equation (4.11), despite its apparent complexity, describes a rich posterior distribution over the values of $f$ that is analytically computable and whose design depends essentially only on the definition of the kernel map (4.10).
Examples of kernel functions include the $\mathcal{C}^{\infty}$ squared exponential
$\kappa_{\Sigma}(\lambda,\lambda^{\prime})=\sigma_{0}e^{-r_{\Sigma}^{2}(\lambda,\lambda^{\prime})/2};\qquad r_{\Sigma}^{2}(\lambda,\lambda^{\prime})=\lambda^{\intercal}\Sigma\lambda^{\prime},$
where $\sigma_{0}>0$ is a scale parameter and $\Sigma$ a diagonal matrix with positive length scales $\{\sigma_{i}\}_{i=1}^{m}$ and the Matérn family, which contains kernels (defined as functions of $r_{\Sigma}$, above) with various degree of smoothness [williams2006gaussian]. Kernels intuitively encode the concept of distance or similarity between data points, and allow to sidestep the definition of (possibly obscurer) feature maps and covariance matrices. The series of “free parameters”222 That is, $\sigma_{\varepsilon}$, $\sigma_{0}$ and the diagonal entries of $\Sigma$ when using the squared exponential or a Matérn kernel. that appears in (4.11) – usually called hyperparameters in the Bayesian literature (not to be confused with $\lambda$) – do not have to be fixed constants, but may be estimated from the observations [williams2006gaussian, shahriari:2016:taking_human_out] (automatic relevance determination). Thus, the design choice left to the experimenter mainly boils down the choice of the kernel function.

### v3

@@ 4.1 Model-Based HPO: A General Schema
2:Warm-up dataset $\mathcal{C}_{0}=\{(\lambda_{j},f(\lambda_{j}))\}_{j=1}^{n_{0}}$
@@
[…]
6:Fit $\hat{f}_{0}$ to $\mathcal{C}_{0}$ $\triangleright$ Initialize the surrogate model
@@
8:  $\lambda_{k}\leftarrow\operatorname*{arg\,max}_{\lambda^{\prime}}\alpha(\lambda^{\prime},\hat{f}_{k-1},\beta)$ $\triangleright$ Select next hyperparameter
9:  Evaluate $f(\lambda_{k})$ $\triangleright$ Expensive step
10:  if $f(\lambda_{k})\leq\beta$ then
11:   $\beta\leftarrow f(\lambda_{k})$ $\triangleright$ Update incumbent performance
12:  end if
13:  $\mathcal{C}_{k}\leftarrow\mathcal{C}_{k-1}\cup\{(\lambda_{k},f(\lambda_{k}))\}$
14:  Fit $\hat{f}_{k}$ to $\mathcal{C}_{k}$ $\triangleright$ Update the surrogate model
@@
The vast majority of model-based methods are, in principle, sequential. At each iteration, the surrogate model is used to select the next hyperparameter configuration to evaluate, typically through the optimization of an acquisition function. Then, the response function is evaluated on the selected configuration by running the learning algorithm and assessing the resulting model performance. Finally, the surrogate model is updated with the newly collected data, or refitted on all data collected so far, and the process is repeated.
Algorithm 1 summarizes the typical steps of an MB-HPO method, where the selection of the next configuration is guided by the optimization of an acquisition function. The surrogate model is typically initialized by fitting it to several evaluations obtained during a warm-up stage, where configurations are generated using methods such as random search (see Section 3.2). We denote this initial configuration set as $\mathcal{C}_{0}=\{(\lambda_{j},f(\lambda_{j}))\}$. New configurations are selected and evaluated, then added to the set of points $\mathcal{C}_{k}=\mathcal{C}_{0}\cup\{(\lambda_{i},f(\lambda_{i}))\}_{i=1}^{k}$, where $k\in[T]$ denotes the iteration index and $T\in\mathds{N}^{+}$ is the total number of iterations. This set of points is then used to fit the $k$-th version of $\hat{f}$, which we denote by $\hat{f}_{k}$. We denote the acquisition function as:
$\alpha\mathrel{\mathop{\ordinarycolon}}\Lambda\times\hat{\mathcal{F}}\times\mathcal{K}\to\mathds{R}.$ (4.1)
Here, $\hat{\mathcal{F}}$ denotes the class of surrogate functions and $\mathcal{K}$ optional auxiliary information that may be needed by the acquisition function, including the validation performance of the incumbent — denoted by $\beta$ in the algorithm, so that $\mathcal{K}=\mathds{R}$.
One may interpret the interplay between fitting a surrogate model and choosing the next point to evaluate through an acquisition function as an iterative refinement of a sampling distribution $\nu_{k}\in\mathcal{P}(\Lambda)$ — implicitly defined through $\hat{f}_{k}$ and $\alpha$. In comparison, random and quasi-random search methods discussed in the previous chapter maintain a predetermined distribution $\nu\in\mathcal{P}(\Lambda)$ that does not change through iterations (cf. Section 3.2).
In the next section, we focus on model-based methods rooted in Bayesian optimization. We discuss specific implementations of surrogate models and their update strategies, as well as various acquisition functions.
@@ 4.2 Bayesian Optimization
BO has long been used as a method to globally optimize black-box functions, that is, functions for which we may only observe the value $f(\lambda)$ at any point $\lambda\in\Lambda$. It dates back at least to the 1970s with a series of works by J. Močkus [mockus:1975:bayesian_methods_seeking, mockus1978application]. Although traditionally suited for low- to medium-dimensional problems, BO has proven to be a powerful method for optimizing expensive, non-convex, and highly multi-modal functions. As a result, it has attracted considerable attention for its potential to tackle HPO problems, especially following the publication of three influential papers in the early 2010s [bergstra2011algorithms, hutter_sequential_2011, snoek_practical_2012]. In this section, we give an overview of the general methodology in the HPO context. We refer the reader to the work by jones1998efficient for a more general perspective of the methodology and to the review by shahriari:2016:taking_human_out for further insights into practical aspects and applications to machine learning beyond HPO.
In Bayesian optimization, the surrogate model $\hat{f}$ is a probabilistic model. Specifically, for any configuration $\lambda\in\Lambda$, the surrogate model $\hat{f}(\lambda)$ is a random variable whose distribution captures the uncertainty about the possible values of the true objective function $f(\lambda)$. This uncertainty has an epistemic nature, i.e., due to limited number of observations of configurations-responses pairs. The surrogate is updated using Bayesian inference, conditioning on the collected data $\mathcal{C}_{k}$, to obtain a posterior distribution over functions $\hat{f}_{k}$. This posterior is then used to guide the selection of the next configuration via an acquisition function, as introduced in the previous section. The acquisition function encodes a strategy to balance exploration: sampling in regions of high uncertainty to improve the surrogate model, and exploitation: selecting configurations that are expected to yield good performance. In this context, the surrogate model’s ability to represent epistemic uncertainty is crucial for enabling effective exploration.
Next, we discuss standard choices to design and fit surrogate models in BO, including Gaussian processes and tree Parzen estimators. We discuss acquisition functions, focusing on the expected improvement [mockus1978application]. A classic instantiation of BO consists in modeling the surrogate with a Gaussian process and using expected improvement as the acquisition function. It is well-suited for the global optimization of smooth multivariate real functions, and it provides an accessible setting to implement BO for hyperparameter optimization.
@@ 4.2.1 Bayesian Linear Model
To introduce our first surrogate model, we start with a simple linear model $h_{w}$ in a given feature space $\mathcal{V}\subseteq\mathds{R}^{d}$, to which we add independent Gaussian noise with variance $\sigma^{2}$:
@@
where $\varphi\mathrel{\mathop{\ordinarycolon}}\Lambda\to\mathcal{V}$ is a feature map and $w\in\mathds{R}^{d}$ is a weight vector. Using linear maps directly in the input space would result in an overly simplistic model — we are trying to approximate a response function that can be highly non-convex and multimodal. However, mapping configurations into a suitable feature space allows us to capture much richer behaviors while retaining many of the benefits of the linear formulation [murphy2012machine, Ch. 14]. As evaluating $f$ is expensive, we expect to work with only a few observations. The noise in Eq. (4.2) models potential errors (and inherent stochasticity) in the observations (we further discuss this aspect in Section 4.2.4).
Suppose now that we are at the $k$-th iteration of the algorithm, after having collected the observations $\mathcal{C}_{k}=\{(\lambda_{i},f(\lambda_{i}))\}_{i=1}^{k}$.  [^1] In the following, we disregard the warm-up set, for simplicity. Let
$Y_{k}=\left(f(\lambda_{1}),\dots,f(\lambda_{k})\right)^{\intercal}\in\mathds{R}^{k}$
be the vector of observed values, and denote by
$X_{k}=\left(\varphi(\lambda_{1}),\dots,\varphi(\lambda_{k})\right)\in\mathds{R}^{d\times k}$
@@
the corresponding design matrix in the feature space (organized by columns). We are interested in obtaining a probability distribution over the output of $f$ at a query point $\lambda_{*}\in\Lambda$, that is, in computing
$y_{*}\sim\hat{f}_{k}(\lambda_{*})=p(\cdot\mid x_{*},X_{k},Y_{k}),$
where $x_{*}=\varphi(\lambda_{*})$ denotes the feature representation of the query point $\lambda_{*}$. Here, the surrogate model $\hat{f}_{k}\mathrel{\mathop{\ordinarycolon}}\Lambda\to\mathcal{P}(\mathds{R})$ produces a predictive distribution over real-valued outcomes, modeling the response function $\lambda$. The resulting distribution will be employed in Section 4.2.3, where the acquisition function uses the surrogate $\hat{f}_{k}$ to guide the selection of the next configuration $\lambda_{k+1}$.
Modeling uncertainty away from observations.
@@
We could follow an empirical risk minimization approach, obtaining the weights $\bar{w}_{k}\in\mathds{R}^{d}$ of the model in Eq. (4.2) by minimizing, for example, a (regularized) mean squared error over the dataset $(X_{k},Y_{k})$. We could then directly set $\hat{f}_{k}=h_{\bar{w}_{k}}$. Such an estimate would, however, leave us with a rather meager predictive distribution $p(\cdot\mid x_{*},X_{k},Y_{k},\bar{w}_{k})$, capable of capturing only our belief of the amount of noise present in the observations. As we discussed above, we would like the stochasticity of $\hat{f}$ to (primarily) capture our uncertainty about the behavior of $f$ in unexplored regions.
To better understand the difference between these two aspects, following an argument by jones1998efficient, consider the case where $f$ is a (non-linear) smooth deterministic function. If we were to follow an empirical risk minimization (ERM) approach, we could either omit the noise term $\varepsilon$ in Eq. (4.2), or leave it for capturing modeling errors. However, the first choice would yield a single point estimate for any $\lambda\in\Lambda$, which is very unlikely to be accurate, especially for points far from those previously observed. Conversely, if we still wish to maintain a noise component, $\hat{f}_{k}$ would trivially be inaccurate at $\lambda_{i}$ for $i\in[k]$ (since $f$ is deterministic). Thus, both ERM modeling options would leave us with a surrogate model that is almost surely inaccurate.
@@
A way to escape this apparent stalemate is to replace the independent Gaussian noise of Eq. (4.2) with one that depends on the evaluation point, i.e. letting $\varepsilon\mathrel{\mathop{\ordinarycolon}}\Lambda\to\mathds{R}$ be a function of $\lambda$. When fitting the model to the $k$ observations, this choice allows having $\varepsilon_{k}(\lambda_{i})=0$ for $i\in[k]$ while still retaining $\varepsilon_{k}(\lambda)\neq 0$ for $\lambda\neq\lambda_{i}$. Furthermore, since we assumed $f$ to be smooth, $\varepsilon_{k}$ will also be smooth (as a difference of smooth functions). By continuity, we then have $\varepsilon_{k}(\lambda)\to\varepsilon_{k}(\lambda_{i})=0$ as $\lambda\to\lambda_{i}$. The Bayesian perspective offers a principled way to achieve this behavior, as we discuss next.
Deriving the posterior predictive distribution.
We can reason about all linear models in the feature space that are consistent with the observations, weighted by their posterior probability. In other words, we construct the surrogate model as a weighted average of all plausible $h_{w}$ given the observed data. Formally, we marginalize over the weights $w$ and set
$\hat{f}_{k}(\lambda_{*})=\int p(\cdot\mid x_{*},X_{k},Y_{k},w)p(w\mid X_{k},Y_{k})\,\mathrm{d}w\in\mathcal{P}(\mathds{R}).$ (4.3)
@@
Given our modeling assumption of Eq. (4.2) and the hypothesis that the noise is independent, the first distribution simply reads as
$p(\cdot\mid x_{*},X_{k},Y_{k},w)=p(\cdot\mid x_{*},w)=\mathcal{N}(h_{w}(\lambda_{*}),\sigma_{\varepsilon}^{2})=\mathcal{N}(x_{*}^{\intercal}w,\sigma_{\varepsilon}^{2}).$ (4.4)
By the Bayes’ rule, $p(\cdot\mid X_{k},Y_{k})$ is instead given by
$p(\cdot\mid X_{k},Y_{k})=\frac{p_{y}(\cdot\mid X_{k},w)p(\cdot\mid X_{k})}{p_{y}(\cdot\mid X_{k})}.$ (4.5)
@@
$p_{y}(\cdot\mid X_{k},w)=\mathcal{N}(X_{k}^{\intercal}w,\sigma_{\varepsilon}^{2}I).$
@@
$p(\cdot\mid X_{k})=p=\mathcal{N}(0,\Sigma),$ (4.6)
where $\Sigma$ is a positive definite covariance matrix. This is a so-called prior. Equipped with Eq. (4.6), we can now also compute the denominator of Eq. (4.5) by marginalizing $p_{y}(\cdot\mid X_{k})$ over $w$. We obtain
$p_{y}(\cdot\mid X_{k})=\int p_{y}(\cdot\mid X_{k},w)p(w)\,\mathrm{d}w=\mathcal{N}(0,X_{k}^{\intercal}\Sigma X_{k}+\sigma_{\varepsilon}^{2}I).$ (4.7)
We recognize, now, that the feature map only appears in a quadratic form where the $(i,j)$-th entry is given by $\varphi(\lambda_{i})^{\intercal}\Sigma\,\varphi(\lambda_{j})$: a dot product in the feature space. We then introduce the kernel corresponding to the feature map $\varphi$ as follows:
@@
and rewrite Eq. (4.7) in terms of the Gram matrix $K=\{\kappa_{\Sigma}(\lambda_{i},\lambda_{j})\}_{i,j=1}^{k}\in\mathds{R}^{k\times k}$ as
$p_{y}(\cdot\mid X_{k})=\mathcal{N}(0,K+\sigma_{\varepsilon}^{2}I).$
Finally, we have all the terms to compute Eq. (4.3); which is, again, Gaussian:
$\hat{f}_{k}(\lambda_{*})=p(\cdot\mid x_{*},X_{k},Y_{k})=\\ \mathcal{N}(k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)^{-1}Y_{k},\kappa_{\Sigma}(\lambda_{*},\lambda_{*})-k_{*}^{\intercal}(K+\sigma_{\varepsilon}^{2}I)^{-1}k_{*})$ (4.9)
where $k_{*}=\{\kappa_{\Sigma}(\lambda_{*},\lambda_{i})\}_{i=1}^{k}\in\mathds{R}^{k}$ is the vector of kernel evaluations between the query and the observation points. Eq. (4.9), despite its apparent complexity, describes a rich posterior distribution over the values of $f$ that is analytically computable and whose design depends essentially only on the definition of the kernel map in Eq. (4.8).
An example of kernel function is the squared exponential kernel:
$\kappa_{\Sigma}(\lambda,\lambda^{\prime})=\sigma_{0}e^{-r_{\Sigma}^{2}(\lambda,\lambda^{\prime})/2};\qquad r_{\Sigma}^{2}(\lambda,\lambda^{\prime})=(\lambda-\lambda^{\prime})^{\intercal}\Sigma^{-1}(\lambda-\lambda^{\prime}),$
where $\sigma_{0}>0$ is a scale parameter and $\Sigma$ a diagonal matrix with positive length scales $\{\sigma_{i}\}_{i=1}^{m}$. Other examples include kernels from Matérn family, which contains mappings with various degrees of smoothness, defined as functions of $r_{\Sigma}$, introduced above [williams2006gaussian]. Kernels intuitively encode the concept of distance or similarity between data points, and allow sidestepping the definition of explicit feature maps and covariance matrices. The series of “free parameters” that appears in Eq. (4.9) — that is, $\sigma_{\varepsilon}$, $\sigma_{0}$ and the diagonal entries of $\Sigma$ when using the squared exponential or a Matérn kernel — are usually called hyperparameters in the Bayesian literature, and are not to be confused with $\lambda$. They do not need to be fixed constants but can instead be estimated from the observed data — see automatic relevance determination [mackay1992bayesian, williams2006gaussian]. Thus, the design choice left to the experimenter mainly boils down to the choice of the kernel function.